# LexGenius: An Expert-Level Benchmark for Large Language Models in Legal General Intelligence

### v1

@@ Appendix B Definitions of Legal Intelligence Abilities
This section provides detailed definitions for the 20 atomic legal intelligence abilities in the LexGenius framework, following the ability names used in Figure 2. Each ability represents a measurable unit of legal general intelligence, assessed through standardized multiple-choice questions.
@@
Appendix C Experimental setup details
C.1 Large Language Models
@@
We evaluated twelve popular and SOTA LLMs on the proposed LexGenius. For GPT-4o mini, GPT-4 nano, DeepSeek V3, and DeepSeek R1, we accessed the models via their official APIs. For other LLMs, we conducted experimental tests using the official weights. The twelve popular SOTA LLMs are as follows:
@@
Qwen2.5-1.5B-Instruct Hui et al. (2024): A lightweight instruction-tuned model released by Alibaba with 1.5B parameters, designed for edge deployment and local inference with bilingual support and basic task execution.
@@
Qwen2.5-7B-Instruct Hui et al. (2024): A mid-scale model in the Qwen2.5 series, optimized for stronger reasoning and instruction following, suitable for more complex language tasks in medium-sized deployments.
@@
Qwen3-4B Yang et al. (2025): A 4B-parameter model from the third-generation Qwen series, showing strong performance in multilingual, coding, and logical tasks.
@@
Qwen3-8B Yang et al. (2025): An enhanced version of Qwen3 with extended context length and multilingual capabilities, significantly improving performance in complex reasoning and generation tasks.
@@
LLaMA-3.2-1B-Instruct Grattafiori et al. (2024): A compact instruction-tuned model from Meta’s LLaMA 3 series, designed for resource-constrained environments while maintaining core instruction-following capabilities.
@@
LLaMA-3.2-8B-Instruct Grattafiori et al. (2024): A standard model in the LLaMA 3 lineup, offering high-quality multilingual understanding, code generation, and reasoning, achieving state-of-the-art results across many tasks.
@@
GLM-4-9B-Chat GLM et al. (2024): A 9B bilingual chat model developed by Zhipu AI, equipped for multi-turn dialogue, tool use, and contextual memory, with strong performance particularly in Chinese semantic understanding.
@@
DeepSeek-LLM-7B-Chat Bi et al. (2024): A 7B bilingual chat model by DeepSeek, integrating capabilities in code generation, mathematics, and language understanding, suitable for dialogue and multitask settings.
@@
DeepSeek-R1 Guo et al. (2025): A large-scale open-source language model developed by the Chinese company DeepSeek, featuring strong capabilities in mathematics, programming, and reasoning with efficient training and leading performance.
@@
DeepSeek-V3 Liu et al. (2024): A large language model based on a mixture-of-experts architecture, excelling in mathematics, programming, and logical reasoning, and is well-suited for a variety of intelligent application scenarios.
@@
GPT-4o mini Hurst et al. (2024): A parameter-efficient version of OpenAI’s GPT-4o model, supporting multimodal inputs (text, image, audio) with consistent alignment behavior as its full-size counterpart.
@@
GPT-4.1 nano Hurst et al. (2024): An ultra-compact model in the GPT-4.1 series, designed for on-device and embedded inference with support for moderate-length contexts and basic reasoning under resource constraints.
@@
C.2 Two prompt methods
We employed two prompting strategies to evaluate the legal general intelligence on the LexGenius. One is a simple prompt (naive), and the other is a Chain-of-Thought (CoT) prompt. These two strategies are illustrated in Figure 29.
Naive Prompt is the most basic form of prompt design, where the task objective is directly stated to the language model, with the expectation that it will generate the final output without revealing any intermediate reasoning steps. This approach relies on the model’s end-to-end learning ability, requiring it to complete the full input-to-output mapping within a single forward pass. Its main advantage lies in simplicity and speed, but it places high demands on the model’s implicit knowledge and generalization capability. Especially for complex tasks, it is prone to errors due to semantic ambiguity or interference from multi-step reasoning chains. Conceptually, this prompting strategy does not explicitly activate the model’s thinking path, which may lead it to overlook latent contextual connections, implicit assumptions, or task structure. As task complexity increases, performance tends to degrade rapidly due to the lack of a mechanism for step-by-step reasoning construction.
CoT Prompt is to explicitly guide the model through step-by-step reasoning by decomposing a task into a sequence of intermediate steps expressed in natural language Kojima et al. (2022). This approach encourages the model to simulate the human problem-solving process and functions as a form of prompt programming, activating internal causal chains and logical pathways by embedding intermediate reasoning states within the prompt. Mechanistically, CoT prompting leverages the structured knowledge and implicit logical modeling capabilities that large models naturally acquire during training. When the prompt includes cues such as think, step, or other structured instructions, the model tends to produce coherent reasoning processes rather than skipping directly to the final answer. This reasoning chain helps reduce error rates in complex tasks and improves the model’s ability to handle multi-step dependencies, conditional logic, and task decomposition.
Appendix D Results of 20 Legal Intelligence Abilities
@@
The results of 20 legal intelligence abilities for the 12 SOTA LLMs are as illustrated in Table 7. The LLMs’ legal intelligence abilities decline significantly in tasks requiring deeper abstraction. These tasks involve complex value judgments, cross-domain norm integration, and procedural reasoning—areas where LLMs struggle to match human-like legal cognition. This highlights the need for further model optimization in sociological, ethical, and institutional aspects of legal general intelligence.
Appendix E With Different Enhanced Methods
To accurately compare and evaluate the impact of different optimization and enhancement methods on the legal intelligence capabilities of Large Language Models (LLMs), we selected four LLMs (including Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct, Qwen3-4B, and Qwen3-8B) and experimented with Supervised Fine-Tuning (SFT), Chain-of-Thought (CoT), Retrieval-Augmented Generation (RAG), and Reinforcement Learning (RL) algorithms. We randomly sampled 64 test instances from each of the 20 ability test sets in LexGenius, resulting in 1,280 total samples for evaluation. The remaining 7,105 data samples were used as the training set for SFT and RL, as well as for constructing the retrieval corpus. The experimental results of these LLMs, after applying these enhancement methods across various dimensions and tasks of legal intelligence, are illustrated in Table 8 and Table 9.
Appendix F Correlation Analysis
The average performance of the twelve SOTA LLMs on LexGenius is utilized to conduct correlation analysis on the benchmark (as shown in Figure 30). It illustrates that most of the legal intelligence abilities (left), tasks (upper right), and dimensions (lower right) exhibit low correlations. It shows the effectiveness of the developed questions in our LexGenius, because the low intercorrelation suggests that models cannot rely on general legal heuristics or shallow transfer across domains to perform well; instead, success in one category does not guarantee success in others. This reflects the comprehensive coverage and conceptual independence of our benchmark dimensions, further validating their robustness as an evaluation framework.
Appendix G Limitation and Future Work
To bridge the gap between current benchmark capabilities and the demands of real-world legal practice, we outline several directions for future work. While LexGenius establishes a foundational framework for evaluating legal general intelligence, it remains constrained by unimodal design, limited jurisdictional scope, and a lack of temporal modeling. To address these limitations, future iterations will pursue enhancements in three key areas: incorporating multimodal legal tasks, expanding linguistic and legal diversity, and introducing dynamic temporal reasoning. These improvements aim to more accurately reflect the complexity of legal environments and to enable more robust, generalizable evaluations of large language models. We hope that LexGenius and these discussions can promote the development and research of legal general intelligence.
G.1 Limitation
While LexGenius offers a structured framework for evaluating legal general intelligence in large language models, it still falls short in key areas that limit its realism and generalizability. Specifically, it lacks support for multimodal inputs, is confined to a single language and legal system, and does not account for the temporal dynamics of law. These gaps constrain its alignment with real-world legal complexity and hinder its ability to fully assess models’ applicability in diverse, evolving legal contexts. The following subsections detail these core limitations.
Lack of Multimodal Tasks Limits Realistic Evidence Modeling. The current version of LexGenius is built and evaluated entirely on pure textual materials, without incorporating multimodal evidence types commonly found in real-world cases, such as scanned contract images, video stills, or audio transcriptions. This unimodal design fails to assess a model’s capabilities in visual perception, auditory understanding, and cross-modal reasoning, which are essential for handling actual judicial cases. The absence of multimodal inputs limits the applicability of LLMs in tasks such as evidence review, fact reconstruction, and the interpretation of visual-legal content, thereby reducing the fidelity of the evaluation to real-world legal scenarios.
Linguistic and Jurisdictional Limitations Undermine Cross-Cultural Generalization. LexGenius is currently constructed solely from Chinese-language corpora and based on the civil law system of Mainland China, exhibiting clear singularity in both linguistic and legal dimensions. As a result, model evaluations are only valid within the context of "Chinese civil law" and fail to capture broader capabilities such as interpreting international statutes, translating case law, or conducting comparative legal analysis. In tasks involving global legal services, cross-border disputes, or foreign compliance, this limitation significantly hampers a model’s transferability and generalization, impeding its evolution into a truly universal system of legal intelligence.
Ability Human LLM 1 LLM 2 LLM 3 LLM 4 LLM 5 LLM 6 LLM 7 LLM 8 LLM 9 LLM 10 LLM 11 LLM 12 Naive Prompt Ability 1 89.63 45.00 58.20 50.80 55.20 28.80 37.80 53.00 39.00 64.40 64.80 47.00 41.40 Ability 2 87.41 37.78 60.89 46.22 54.22 30.67 33.33 46.22 34.22 67.11 65.78 47.11 40.44 Ability 3 90.37 49.38 65.15 62.66 65.15 28.63 39.83 59.34 38.59 70.54 70.54 55.19 57.68 Ability 4 92.59 39.40 56.60 48.00 50.60 25.80 32.20 47.60 32.20 59.20 58.00 46.00 39.00 Ability 5 88.89 43.45 55.36 52.38 50.60 23.81 27.98 41.67 35.71 60.71 60.12 39.29 47.62 Ability 6 85.93 40.19 60.29 50.72 55.02 22.49 33.97 54.07 34.93 63.64 64.59 44.50 43.06 Ability 7 85.93 36.20 47.80 41.40 42.80 25.60 31.60 39.40 35.80 46.80 46.80 40.40 35.40 Ability 8 92.59 39.60 52.20 46.00 48.40 23.20 34.80 42.60 27.40 53.00 55.20 40.40 35.80 Ability 9 88.89 31.03 35.78 29.74 25.00 30.60 33.19 24.57 25.43 37.93 37.07 30.17 21.55 Ability 10 86.67 59.00 64.80 58.00 64.20 46.80 56.20 66.00 48.40 67.60 67.60 65.20 59.20 Ability 11 83.70 52.40 63.20 63.80 61.60 42.00 50.80 56.80 41.80 74.00 73.00 59.20 54.40 Ability 12 89.63 43.55 63.55 50.65 60.97 27.74 29.35 58.71 32.26 74.19 73.55 45.81 45.81 Ability 13 85.93 60.40 68.60 64.60 65.20 51.00 63.40 66.20 46.80 68.00 67.40 69.20 60.40 Ability 14 87.41 57.20 71.00 69.00 67.60 45.80 55.00 64.40 51.60 76.80 76.40 67.60 63.00 Ability 15 91.85 56.40 58.00 55.60 54.80 46.00 51.60 57.60 40.20 65.40 66.20 61.20 54.40 Ability 16 82.96 64.40 69.20 62.00 61.20 59.00 59.60 63.60 49.20 75.00 72.60 67.60 59.00 Ability 17 89.63 38.40 45.80 40.80 45.00 23.80 34.20 39.40 33.40 48.80 49.40 40.20 37.00 Ability 18 90.37 50.80 64.60 58.20 66.00 44.00 53.60 60.00 40.40 72.40 71.80 60.40 59.00 Ability 19 79.26 58.20 74.00 65.80 69.60 37.80 54.60 63.80 45.60 80.00 81.20 65.80 61.80 Ability 20 87.41 34.60 45.80 41.40 40.80 26.60 28.80 39.00 29.40 48.40 47.20 42.60 35.00 CoT Prompt Ability 1 89.63 45.00 57.60 50.60 54.40 30.00 36.40 53.20 36.80 64.40 64.20 48.00 56.80 Ability 2 87.41 38.67 60.00 44.44 55.11 28.44 32.00 45.78 30.67 66.67 66.67 46.22 50.22 Ability 3 90.37 48.13 63.90 62.24 66.80 26.56 38.59 60.17 33.20 72.20 71.78 55.19 56.85 Ability 4 92.59 42.00 56.20 48.00 49.20 23.60 32.20 47.00 32.00 60.40 58.80 45.80 40.00 Ability 5 88.89 42.26 54.17 52.98 50.60 28.57 30.95 41.07 27.38 59.52 61.31 40.48 47.62 Ability 6 85.93 42.58 59.33 50.72 58.37 25.36 35.41 54.55 33.49 64.59 64.59 45.45 44.02 Ability 7 85.93 36.00 48.00 42.00 41.40 24.60 32.00 39.60 34.40 46.60 47.60 39.40 39.40 Ability 8 92.59 39.60 51.80 44.20 47.40 22.20 35.00 41.80 29.80 53.60 53.40 40.60 48.00 Ability 9 88.89 31.90 35.78 28.02 22.84 31.90 33.19 28.02 28.45 37.50 38.79 29.31 32.76 Ability 10 86.67 59.40 65.40 58.60 62.80 46.60 55.20 65.60 50.20 68.00 67.40 65.80 60.80 Ability 11 83.70 52.60 61.80 63.40 60.40 40.20 51.00 57.20 36.40 73.20 71.40 58.60 58.40 Ability 12 89.63 44.84 63.23 50.32 60.65 26.45 29.03 59.68 32.58 74.19 73.23 46.45 60.97 Ability 13 85.93 61.80 68.60 64.40 63.20 53.40 64.20 66.80 49.00 68.20 67.80 68.60 64.40 Ability 14 87.41 57.20 70.40 68.20 66.00 41.40 56.40 63.40 50.40 77.00 77.20 67.20 66.20 Ability 15 91.85 56.20 59.60 54.00 54.60 45.60 51.80 57.40 42.60 65.00 65.20 61.40 54.00 Ability 16 82.96 66.00 68.80 61.20 60.80 56.80 61.20 63.60 49.20 75.00 75.40 68.80 61.40 Ability 17 89.63 38.80 45.20 41.20 44.40 21.60 32.40 38.40 30.60 51.20 48.80 39.60 37.00 Ability 18 90.37 50.80 64.60 58.00 65.00 42.80 54.80 60.80 42.60 73.60 72.80 61.80 60.60 Ability 19 79.26 59.20 74.40 66.00 68.40 39.00 56.60 64.60 43.60 79.60 81.00 67.00 66.00 Ability 20 87.41 36.60 46.20 41.80 39.80 25.60 30.40 40.40 29.40 47.00 47.40 42.00 46.80
Lack of Evaluation on Temporal Sensitivity and Legal Validity Awareness. A core characteristic of law is its temporal nature—the applicable rules for a given issue may vary significantly across time, especially before and after legislative amendments. LexGenius currently does not incorporate a systematic temporal dimension to assess whether models can understand the time-bound applicability of statutes, the validity period of precedents, or transitional legal provisions. Without such temporal sensitivity tests, models may produce outdated or legally invalid answers when facing evolving legal frameworks, with no mechanism to detect these errors.
G.2 Future work
LexGenius aims to advance Chinese legal intelligence toward generalization. The current version covers fundamental legal domains and establishes a Dimension–Task–Ability evaluation framework, yet it still falls short of capturing the full complexity of real-world legal systems. Therefore, our future work will primarily focus on the following:
Incorporating Multimodal Tasks to Enhance Realistic Evidence Modeling. The current version is constructed solely on textual materials and does not include multimodal information commonly encountered in real legal cases, such as scanned contracts, courtroom audio transcriptions, or surveillance video stills. The absence of such inputs limits the evaluation of model capabilities in visual perception, auditory comprehension, and cross-modal reasoning, all of which are essential for key legal tasks such as evidence review, fact reconstruction, and interpretation of visual-legal content. In future iterations, we plan to embed images, audio, and other modalities into tasks to assess models’ ability to reason and judge based on heterogeneous, multi-source information, thus aligning evaluation more closely with practical judicial needs.
Expanding Linguistic and Jurisdictional Coverage to Improve Cross-Cultural Generalization. The current dataset is grounded in Chinese-language texts and the civil law system of Mainland China, exhibiting limitations in both language and legal tradition. This restricts the applicability of evaluation outcomes. Future versions will incorporate legal texts from Hong Kong, Macau, and Taiwan, as well as English-language statutes and case law from common law systems. We aim to construct bilingual legal QA pairs, statute translation tasks, and comparative law analyses to support the evaluation of models’ capabilities in understanding, aligning, and adapting across legal and linguistic contexts. This expansion will contribute to the training and benchmarking of legal language models equipped for global legal services.
Introducing Dynamic Testing of Legal Temporality and Time Sensitivity. Legal applicability is highly time-dependent—amendments to laws can lead to drastically different rulings, and precedents often carry specific periods of validity and applicability. Currently, LexGenius lacks a systematic temporal dimension, making it difficult to evaluate whether a model can identify the applicable time windows of statutes, conditions for transitional provisions, or conflicts between old and new laws. Future versions will include temporally structured legal tasks that require models to make dynamic judgments under varying legal timeframes, thereby enhancing their understanding and adaptability to evolving legal systems.

### v2

@@ Appendix B Definitions of Legal Intelligence Abilities
This section provides detailed definitions for the twenty atomic legal intelligence abilities in the proposed LexGenius framework, following the ability names utilized in Figure 2. Each ability represents a measurable unit of legal general intelligence, assessed through standardized multiple-choice questions.
@@
Appendix C Annotation Details
We recruited nine master’s candidates in law for double-blind annotation. As detailed in Section 4.2 and Figure 3, evaluators assessed questions on a 5-point scale based on five specific criteria: statutory accuracy, logical soundness, answer uniqueness, competence alignment, and normative wording. Arbitration was triggered for score discrepancies exceeding two points. Participants were informed of data usage and management policies. To ensure fair labor practices, annotators were compensated at a rate of US\$15 per hour, which exceeds the local average wage for student research assistants.
Appendix D Experimental setup details
D.1 Large Language Models
@@
We evaluated 12 LLMs on LexGenius. For GPT-4o mini, GPT-4 nano, DeepSeek-V3, and DeepSeek-R1, we accessed them via their official APIs. For other LLMs, we conducted experimental tests using the official weights. These LLMs are as follows:
@@
Qwen2.5-1.5B-Instruct hui2024qwen2: A lightweight instruction-tuned model released by Alibaba with 1.5B parameters, designed for edge deployment and local inference with bilingual support and basic task execution.
@@
Qwen2.5-7B-Instruct hui2024qwen2: A mid-scale model in the Qwen2.5 series, optimized for stronger reasoning and instruction following, suitable for more complex language tasks in medium-sized deployments.
@@
Qwen3-4B yang2025qwen3: A 4B-parameter model from the third-generation Qwen series, showing strong performance in multilingual, coding, and logical tasks.
@@
Qwen3-8B yang2025qwen3: An enhanced version of Qwen3 with extended context length and multilingual capabilities, significantly improving performance in reasoning and generation tasks.
@@
LLaMA-3.2-1B-Instruct grattafiori2024llama: A compact instruction-tuned model from Meta’s LLaMA 3 series, designed for resource-constrained environments while maintaining core instruction-following capabilities.
@@
LLaMA-3.2-8B-Instruct grattafiori2024llama: A standard model in the LLaMA 3 lineup, offering high-quality multilingual understanding, code generation, and reasoning, achieving state-of-the-art results across many tasks.
@@
GLM-4-9B-Chat glm2024chatglm: A 9B bilingual chat model developed by Zhipu AI, equipped for multi-turn dialogue, tool use, and contextual memory, with strong performance particularly in Chinese semantic understanding.
@@
DeepSeek-LLM-7B-Chat bi2024deepseek: A 7B bilingual chat model by DeepSeek, integrating capabilities in code generation, mathematics, and language understanding, suitable for dialogue and multitask settings.
@@
DeepSeek-R1 guo2025deepseek: A large-scale open-source language model developed by the Chinese company DeepSeek, featuring strong capabilities in mathematics, programming, and reasoning with efficient training and leading performance.
@@
DeepSeek-V3 liu2024deepseek: A large language model based on a mixture-of-experts architecture, excelling in mathematics, programming, and logical reasoning, and is well-suited for a variety of intelligent application scenarios.
@@
GPT-4o mini hurst2024gpt: A parameter-efficient version of OpenAI’s GPT-4o, supporting multimodal inputs (text, image, audio) with consistent alignment behavior as its full-size counterpart.
@@
GPT-4.1 nano hurst2024gpt: An ultra-compact model in the GPT-4.1 series, designed for on-device and embedded inference with support for moderate-length contexts and basic reasoning under resource constraints.
@@
Ability Human LLM 1 LLM 2 LLM 3 LLM 4 LLM 5 LLM 6 LLM 7 LLM 8 LLM 9 LLM 10 LLM 11 LLM 12
Naive Prompt
Ability 1 89.63 45.00 58.20 50.80 55.20 28.80 37.80 53.00 39.00 64.40 64.80 47.00 41.40
Ability 2 87.41 37.78 60.89 46.22 54.22 30.67 33.33 46.22 34.22 67.11 65.78 47.11 40.44
Ability 3 90.37 49.38 65.15 62.66 65.15 28.63 39.83 59.34 38.59 70.54 70.54 55.19 57.68
Ability 4 92.59 39.40 56.60 48.00 50.60 25.80 32.20 47.60 32.20 59.20 58.00 46.00 39.00
Ability 5 88.89 43.45 55.36 52.38 50.60 23.81 27.98 41.67 35.71 60.71 60.12 39.29 47.62
Ability 6 85.93 40.19 60.29 50.72 55.02 22.49 33.97 54.07 34.93 63.64 64.59 44.50 43.06
Ability 7 85.93 36.20 47.80 41.40 42.80 25.60 31.60 39.40 35.80 46.80 46.80 40.40 35.40
Ability 8 92.59 39.60 52.20 46.00 48.40 23.20 34.80 42.60 27.40 53.00 55.20 40.40 35.80
Ability 9 88.89 31.03 35.78 29.74 25.00 30.60 33.19 24.57 25.43 37.93 37.07 30.17 21.55
Ability 10 86.67 59.00 64.80 58.00 64.20 46.80 56.20 66.00 48.40 67.60 67.60 65.20 59.20
Ability 11 83.70 52.40 63.20 63.80 61.60 42.00 50.80 56.80 41.80 74.00 73.00 59.20 54.40
Ability 12 89.63 43.55 63.55 50.65 60.97 27.74 29.35 58.71 32.26 74.19 73.55 45.81 45.81
Ability 13 85.93 60.40 68.60 64.60 65.20 51.00 63.40 66.20 46.80 68.00 67.40 69.20 60.40
Ability 14 87.41 57.20 71.00 69.00 67.60 45.80 55.00 64.40 51.60 76.80 76.40 67.60 63.00
Ability 15 91.85 56.40 58.00 55.60 54.80 46.00 51.60 57.60 40.20 65.40 66.20 61.20 54.40
Ability 16 82.96 64.40 69.20 62.00 61.20 59.00 59.60 63.60 49.20 75.00 72.60 67.60 59.00
Ability 17 89.63 38.40 45.80 40.80 45.00 23.80 34.20 39.40 33.40 48.80 49.40 40.20 37.00
Ability 18 90.37 50.80 64.60 58.20 66.00 44.00 53.60 60.00 40.40 72.40 71.80 60.40 59.00
Ability 19 79.26 58.20 74.00 65.80 69.60 37.80 54.60 63.80 45.60 80.00 81.20 65.80 61.80
Ability 20 87.41 34.60 45.80 41.40 40.80 26.60 28.80 39.00 29.40 48.40 47.20 42.60 35.00
CoT Prompt
Ability 1 89.63 45.00 57.60 50.60 54.40 30.00 36.40 53.20 36.80 64.40 64.20 48.00 56.80
Ability 2 87.41 38.67 60.00 44.44 55.11 28.44 32.00 45.78 30.67 66.67 66.67 46.22 50.22
Ability 3 90.37 48.13 63.90 62.24 66.80 26.56 38.59 60.17 33.20 72.20 71.78 55.19 56.85
Ability 4 92.59 42.00 56.20 48.00 49.20 23.60 32.20 47.00 32.00 60.40 58.80 45.80 40.00
Ability 5 88.89 42.26 54.17 52.98 50.60 28.57 30.95 41.07 27.38 59.52 61.31 40.48 47.62
Ability 6 85.93 42.58 59.33 50.72 58.37 25.36 35.41 54.55 33.49 64.59 64.59 45.45 44.02
Ability 7 85.93 36.00 48.00 42.00 41.40 24.60 32.00 39.60 34.40 46.60 47.60 39.40 39.40
Ability 8 92.59 39.60 51.80 44.20 47.40 22.20 35.00 41.80 29.80 53.60 53.40 40.60 48.00
Ability 9 88.89 31.90 35.78 28.02 22.84 31.90 33.19 28.02 28.45 37.50 38.79 29.31 32.76
Ability 10 86.67 59.40 65.40 58.60 62.80 46.60 55.20 65.60 50.20 68.00 67.40 65.80 60.80
Ability 11 83.70 52.60 61.80 63.40 60.40 40.20 51.00 57.20 36.40 73.20 71.40 58.60 58.40
Ability 12 89.63 44.84 63.23 50.32 60.65 26.45 29.03 59.68 32.58 74.19 73.23 46.45 60.97
Ability 13 85.93 61.80 68.60 64.40 63.20 53.40 64.20 66.80 49.00 68.20 67.80 68.60 64.40
Ability 14 87.41 57.20 70.40 68.20 66.00 41.40 56.40 63.40 50.40 77.00 77.20 67.20 66.20
Ability 15 91.85 56.20 59.60 54.00 54.60 45.60 51.80 57.40 42.60 65.00 65.20 61.40 54.00
Ability 16 82.96 66.00 68.80 61.20 60.80 56.80 61.20 63.60 49.20 75.00 75.40 68.80 61.40
Ability 17 89.63 38.80 45.20 41.20 44.40 21.60 32.40 38.40 30.60 51.20 48.80 39.60 37.00
Ability 18 90.37 50.80 64.60 58.00 65.00 42.80 54.80 60.80 42.60 73.60 72.80 61.80 60.60
Ability 19 79.26 59.20 74.40 66.00 68.40 39.00 56.60 64.60 43.60 79.60 81.00 67.00 66.00
Ability 20 87.41 36.60 46.20 41.80 39.80 25.60 30.40 40.40 29.40 47.00 47.40 42.00 46.80
D.2 Two Prompt Methods
To evaluate LexGenius, we employed Naive and CoT strategies (see Figure 29). The Naive prompt prioritizes efficiency via direct output generated in a single pass, though its lack of reasoning often degrades performance on complex tasks. Conversely, CoT simulates human problem-solving by decomposing tasks into intermediate steps, activating causal chains to significantly reduce errors in multi-step dependencies and conditional logic.
Appendix E Results of Twenty Abilities
@@
The results of 20 legal intelligence abilities for the 12 LLMs are shown in Table 7. The LLMs’ legal intelligence abilities decline significantly in tasks requiring deeper abstraction. These tasks involve complex value judgments, cross-domain norm integration, and procedural reasoning—areas where LLMs struggle to match human-like legal cognition. This highlights the need for further model optimization in sociological, ethical, and institutional aspects of legal general intelligence.
@@
Appendix F With Different Enhanced Methods
To evaluate the impact of different optimization and enhancement methods on the legal intelligence capabilities of LLMs, we selected four LLMs (including Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct, Qwen3-4B, and Qwen3-8B) and experimented with Supervised Fine-Tuning (SFT), Chain-of-Thought (CoT), Retrieval-Augmented Generation (RAG), and Reinforcement Learning (RL) algorithms. We randomly sampled 64 test instances from each of the 20 ability test sets in LexGenius, resulting in 1,280 total samples for evaluation. The remaining 7,105 data samples were used as the training set for SFT and RL, as well as for constructing the retrieval corpus. The appropriate parameters were selected for SFT and RL training. The experimental results of these LLMs, after applying these enhancement methods across various dimensions and tasks of legal intelligence, are shown in Table 8 and Table 9.
Appendix G Correlation Analysis
The performance of 12 LLMs on LexGenius is utilized to analyze correlations (see Figure 30). It illustrates that most of the legal intelligence abilities (left), tasks (upper right), and dimensions (lower right) exhibit low correlations. It shows the effectiveness of LexGenius because the low intercorrelation suggests LLMs cannot rely on general legal heuristics or shallow transfer across domains to perform well; instead, success in one category does not guarantee success in others. This reflects the comprehensive coverage and conceptual independence of our benchmark dimensions, further validating their robustness as an evaluation framework.
Appendix H Limitations
Although LexGenius structures legal general intelligence evaluation, it is limited by a lack of multimodal capabilities, cross-jurisdictional coverage, and temporal awareness. These gaps constrain its ability to capture real-world complexity. The following subsections detail these core limitations.
Lack of Multimodal Tasks Limits Realistic Evidence Modeling. The current version of LexGenius relies entirely on pure textual materials, excluding multimodal evidence types common in real-world cases, such as scanned contracts, video stills, or audio transcriptions. This unimodal design fails to assess capabilities in visual perception, auditory understanding, and cross-modal reasoning essential for handling actual judicial cases. Consequently, the absence of multimodal inputs limits LLM applicability in tasks such as evidence review, fact reconstruction, and visual-legal interpretation, reducing evaluation fidelity to real-world scenarios.
Linguistic and Jurisdictional Limitations Undermine Cross-Cultural Generalization. LexGenius is currently constructed solely from Chinese corpora and Mainland China’s law system, exhibiting distinct linguistic and legal singularity. Consequently, evaluations are confined to this context, failing to capture broader capabilities like interpreting international statutes or comparative analysis. This restriction limits applicability in global legal services and cross-border disputes, hampering transferability and impeding the evolution into a universal legal intelligence system.
Lack of Evaluation on Temporal Sensitivity and Legal Validity Awareness. A core characteristic of law is its temporal nature. Applicable rules for a given issue may vary across time, especially before and after legislative amendments. LexGenius currently does not incorporate a systematic temporal dimension to assess whether models can understand the time-bound applicability of statutes, the validity period of precedents, or transitional legal provisions. Without such temporal sensitivity tests, models may produce outdated or legally invalid answers when facing evolving legal frameworks, with no mechanism to detect these errors.
Appendix I Future Work
While LexGenius establishes a structured evaluation framework for Chinese legal general intelligence, it has yet to fully capture real-world complexity. Therefore, our future work focuses on:
Incorporating Multimodal Tasks to Enhance Realistic Evidence Modeling. The current version of LexGenius relies solely on text and does not include multimodal information common in real legal cases, such as scanned contracts, courtroom audio, or surveillance stills. The absence of such inputs limits the evaluation of model capabilities in visual perception, auditory comprehension, and cross-modal reasoning, essential for evidence review, fact reconstruction, and interpretation of visual-legal content. In the future, we plan to embed images, audio, and other modalities into tasks to assess reasoning capabilities based on heterogeneous, multi-source information, thus aligning evaluation more closely with practical judicial needs.
Expanding Linguistic and Jurisdictional Coverage to Improve Cross-Cultural Generalization. The current dataset is grounded in Chinese texts and the law system of Mainland China, exhibiting limitations in language and legal tradition. This restricts evaluation applicability. Future versions will incorporate texts from Hong Kong, Macau, and Taiwan, as well as English statutes and case law from common law systems. We aim to construct bilingual QA pairs, translation tasks, and comparative analyses to evaluate models’ capabilities in understanding, aligning, and adapting across legal and linguistic contexts. This expansion contributes to benchmarking models for global legal services.
Introducing Dynamic Testing of Legal Temporality and Time Sensitivity. Legal applicability is highly time-dependent. Legal amendments can lead to different rulings, and precedents often carry specific periods of validity and applicability. Currently, LexGenius lacks a systematic temporal dimension, making it difficult to evaluate whether a model can identify the applicable time windows of statutes, conditions for transitional provisions, or conflicts between old and new laws. Future versions will include temporally structured legal tasks that require models to make dynamic judgments under varying timeframes, enhancing their understanding and adaptability to evolving legal systems.